OpenStudy (anonymous):

Find the slope of the sides of the following quadrilaterals. R(-4, 0), S(-2, 6) , T(4, 4), U(2, -2) mRS = mST = mTU = mUR =

OpenStudy (binary3i):

use the same slope formula. p1(x1,y1) and p2(x2,y2) slope =(y2-y1)/(x2-x1)

OpenStudy (campbell_st):

the gradient/slope formula is \[m = \frac{y _{2}-y _{1}}{x _{2}-x _{1}}\] them mRS = (6 - 0)/(-2 + 4) mRS = 3 same process for ST, TU adn UR